40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Jul 8 20:50:49 2011, sonali, Update, Green Locking, Power of the AUX laser increased. Graph3.pngETMY_beam_powers.png
    Reply  Thu Jul 14 02:32:11 2011, sonali, Update, Green Locking, Power of the AUX laser increased. 
       Reply  Fri Jul 15 13:48:56 2011, sonali, Update, Green Locking, Power of the AUX laser increased. ETMY_after_fibre_coupling_labelled.pdf
Message ID: 4958     Entry time: Fri Jul 8 20:50:49 2011     Reply to this: 4965
Author: sonali 
Type: Update 
Category: Green Locking 
Subject: Power of the AUX laser increased. 

The ETMY laser was operating at 1.5 A current and 197 mW power.

For the efficient frequency doubling of the AUX  laser beam at the ETMY table, a higher power is required.

Steve and I changed the current level of the laser from 1.5 A to 2.1 A in steps of 0.1 A and noted the corresponding power output . The graph is attached here.

The laser has been set to current 1.8 Amperes. At this current, the power of the output beam just near the laser output is measured to be 390 mW.

The power of the beam which is being coupled into the optical fibre is measured to be between 159 mW to 164 mW (The power meter was showing fluctuating readings).

The power out of the beam coming out of the fibre far-end at the PSL table is measured to be 72 mW. Here, I have attached a picture of the beam paths of the ETMY table with the beams labelled with their respective powers.

Next we are going to adjust the green alignment on the ETMY and then measure the power of the beam.

At the output end of the fibre on the PSL, a power meter has been put to dump the beam for now as well as to help with the alignment at the ETMY table.

Attachment 1: Graph3.png  610 kB  | Hide | Hide all
Graph3.png
Attachment 2: ETMY_beam_powers.png  324 kB  | Hide | Hide all
ETMY_beam_powers.png
ELOG V3.1.3-